### v1

# Impedance-Engineered Josephson Parametric Amplifier with Single-Step Lithography

Impedance-Engineered Josephson Parametric Amplifier with Single-Step Lithography
Lipi Patel Department of Instrumentation and Applied Physics, Indian Institute of Science, India lipipatel@iisc.ac.in    Samarth Hawaldar Department of Instrumentation and Applied Physics, Indian Institute of Science, India Institute of Science and Technology Austria, 3400 Klosterneuburg, Austria    Aditya Panikkar Department of Instrumentation and Applied Physics, Indian Institute of Science, India    Athreya Shankar Department of Physics, Indian Institute of Technology Madras, Chennai 600036, India Center for Quantum Information, Communication and Computing, Indian Institute of Technology Madras, Chennai 600036, India    Baladitya Suri Department of Instrumentation and Applied Physics, Indian Institute of Science, India
(12 luglio 2025)
Sommario
We present the experimental demonstration of an impedance-engineered Josephson parametric amplifier (IEJPA) fabricated in a single-step lithography process. Impedance engineering is implemented using a lumped-element series LC circuit. We use a simpler lithography process where the entire device – impedance transformer and JPA – are patterned in a single electron beam lithography step, followed by a double-angle Dolan bridge technique for Al-AlOx-Al deposition. We observe nearly quantum-limited amplification with 18 dB gain over a wide 400400400\,400MHz bandwidth centered around 5.35.35.3\,5.3GHz, and a saturation power of $-114\,$dBm. To accurately explain our experimental results, we extend existing theories for impedance-engineered JPAs to incorporate the full sine nonlinearity of both the JPA and the transformer. Our work shows a path to simpler realization of broadband JPAs and provides a theoretical foundation for a novel regime of JPA operation.
Josephson Parametric Amplifiers (JPAs) Yurke et al. (1989) have been a mainstay of superconducting circuit QED systems Vijay, Slichter, and Siddiqi (2011); Hatridge et al. (2013); Lin et al. (2013); Roy and Devoret (2016) due to their ability to provide quantum-limited amplificationCaves (1982), facilitate single-shot readoutKrantz et al. (2016), generate squeezed states of lightMallet et al. (2011); Yurke et al. (1988), and to generate two-mode squeezed states through three-wave and four-wave mixing mechanismsEichler et al. (2011). The phase-insensitive amplification mode of these amplifiers is usually characterized by their gain-bandwidth product defined by the maximum gain times the full width at half maximum of the amplifier. Conventional JPAs are usually narrow-band amplifiers with low gain-bandwidth products (202020\,20dB gain over $20-50\,$MHz) Castellanos-Beltran and Lehnert (2007); Yamamoto et al. (2008); Mutus et al. (2013); Zhou et al. (2014), which limit their utility to amplifying a signal composed of a single mode/tone (frequency component). On the other hand, amplifiers with higher gain-bandwidth products can be used to amplify a range of frequencies over a wider bandwidth. This finds applications in multiplexing qubit readout by allowing the amplification of multiple readout tones within the broad bandwidth of the JPA Vijay (2020).
In order to increase the bandwidth without compromising on the gain, and thereby boost the gain-bandwidth product, Travelling Wave Parametric Amplifiers (TWPAs) CULLEN (1958); Feldman, Parrish, and Chiao (1975); Wahlsten, Rudner, and Claeson (1977); Ho Eom et al. (2012); Yaakobi et al. (2013); Zorin (2019) were developed. TWPAs are transmission lines with a distributed nonlinear inductance that typically provide amplification of around 202020\,20dB over a 444\,4GHz bandwidth. However, they involve complex fabrication techniques that are required to create thousands of identical unit cells with Josephson junctions (JJs) and capacitors Macklin et al. (2015); White et al. (2015).
As simpler alternatives to the TWPA, two important techniques have been proposed and implemented to enhance the gain-bandwidth product of JPAs – impedance-matching and impedance-engineering. Impedance-matching involves coupling the JPA oscillator, consisting of a single rf-SQUID and a shunting capacitor, to the external environment via an impedance transformer like a Klopfenstein taper D.M.Pozar to match the impedance of the parametric oscillator (which is of the order of a few Ohms) to the $50\Omega$ environment. However, this method also requires complex fabrication techniques to create hybrid microstrip-coplanar waveguide (CPW) transmission lines, with a large centimetre-scale footprint Lu et al. (2022); Mutus et al. (2014). The other method of increasing bandwidth, impedance engineeringRoy et al. (2015), involves modifying the environment seen by the JPA by introducing a transformer via a $\lambda/2$ CPW section Roy et al. (2015); Grebel et al. (2021) or a series LC resonator Roy et al. (2015). Conventionally, these devices are fabricated in a multi-step lithography process, that uses a sequence of photo and e-beam lithography steps, interspersed by dielectric and metal deposition steps.
In this Letter, we demonstrate an impedance-engineered JPA optimized for four-wave mixing-based amplification that is fabricated using a single-step lithography technique. The simplicity of our process, in turn, reduces the cost and time needed to fabricate the device. It also reduces the errors in fabrication and thereby increases the reliability and yield of the devices. The device used in this work has a simple design made up entirely of coplanar elements. Operating the device in reflection mode, we demonstrate a high gain over a wide bandwidth with almost quantum-limited noise performance, comparable to state-of-the-art devices Roy et al. (2015); Grebel et al. (2021); Lu et al. (2022). Furthermore, to explain our results, we develop a theoretical model to compute the expected gain of an impedance-engineered JPA accounting for the full non-linearity of both the JPA and the impedance transformer. Our model is a substantial extension of previous theoretical work where only the lowest order non-linearity of the JPA is considered and the transformer is assumed to be made of a linear inductor.
Figura 1: False-colored optical micrograph of an Impedance-engineered JPA. (a) Micrograph of the full IEJPA. From the left to right, the green colored component is the input pad, the blue colored components make up the impedance transformer, and the yellow colored elements make up the JPA. (b) Zoomed-in micrograph of the JJ array of the transformer. (c) Zoomed-in micrograph of the SQUID in the JPA.
Figure 1 shows the false color micrograph of the device and the zoomed-in images of an array of JJs of the transformer and the JPA SQUID. The JPA is formed by a dc-SQUID (nominal area of $138\,\mu$m2 and unbiased critical current of $1.8\,\mu$A) shunting an interdigital capacitor (IDC) with a simulated capacitance $C_{J}\approx 1.6\,$pF to ground. The impedance transformer, comprising an array of three nominally identical Josephson junctions, each with a critical current of 585585585\,585nA, connected in series with another IDC (with a simulated capacitance of $C_{T}\approx 420\,$fF), was placed in series between the transmission line and JPA, forming a galvanic connection with both structures. The JPA and transformer are shown in the false color micrograph in yellow and blue color respectively. To design the device layout, we used Qiskit Metal and we performed eigenmode simulations of the JPA and the impedance transformer using ANSYS HFSS. For estimating the capacitance of the JPA and the impedance transformer, we used ANSYS HFSS along with COMSOL MULTIPHYSICS. Using these simulations, in addition to the capacitances quoted above, we estimate the geometric inductance for the JPA and the transformer to be 0.120.120.12\,0.12nH and 0.200.200.20\,0.20nH, respectively.
The device was fabricated on an intrinsic silicon substrate using an aluminum evaporative deposition process. First, we did a preliminary cleaning of the substrate using “piranha” (a $3:1$ volume-ratio mixture of $\text{H}_{2}\text{SO}_{4}$ and $\text{H}_{2}\text{O}_{2}$, for 10 minutes) followed by dilute hydrofluoric acid(HF) – a $50:1$ volume-ratio mixture of DI water and HF, for 30 seconds. Next, a resist bi-layer comprising a 500500500\,500nm thick bottom layer of methacrylic acid (MAA) in methyl methacrylate (MMA), diluted in ethyl lactate (EL) at a 11% concentration, and a 200200200\,200nm thick top layer of 4% PMMA 950K in Anisole was spin-coated on the substrate. We then carried out e-beam lithography in a Raith e-Line at 202020\,20kV acceleration voltage. We pattern both the impedance transformer and the JPA in this step. The pattern development was done using the MIBK (Methyl Isobutyl Ketone) developer. Subsequently, we used the Dolan-bridge technique for forming the Josephson junctions through the double-angle (shadow) evaporation of aluminum. The double-angle deposition was done at ±$\pm 45^{\circ}$ with the first and second layers of Al having thicknesses 202020\,20nm and 404040\,40nm, respectively. All the Josephson junctions in the circuit – both in the JPA SQUID loop and the impedance transformer – were formed in a single in-situ oxidation step between the double-angle Al depositions. The oxidation was done at 700 mTorr pressure for 20 minutes.
The device was packaged inside a copper sample box using gold wire bonds. The sample box was then mounted on the mixing chamber plate of a dilution refrigerator at 151515\,15mK. To tune the frequency of the JPA, we applied a DC magnetic flux using an external coil made of superconducting wire. The sample box containing the JPA device was mounted at the center of the external coil such that the plane of the SQUID of the JPA was perpendicular to the direction of the magnetic field of the coil. The resonance frequency $\Omega_{J}$ of the JPA could then be varied using the external magnetic flux $\Phi_{\text{DC}}$. To further prevent noise due to external magnetic fields, the sample box, along with the magnetic coil, was placed inside a superconducting aluminum shield mounted at the mixing chamber.
[…]
The noise photon added of the IEJPA ($n_{\text{add}}$) is estimated by making use of the SNR extracted from the spectrum measured using a spectrum analyser in the “pump on” and “pump off” cases, and using the Friis formula for noise. In summary, if we assume that the high electron-mobility transistor (HEMT) amplifier has a high enough gain and added noise such that the effect of noise added by the amplifiers further in the chain can be ignored, the noise photons added, $n_{\text{add}}$ by the IEJPA can be estimated using
$n_{\text{add}}=T_{\text{HEMT}}\Bigl{[}\frac{\text{SNR}_{\text{off}}}{\text{SNR}_{\text{on}}}-\frac{L}{G}\Bigr{]}\frac{k_{b}}{\hbar\omega},$ (1)
where $n_{\text{add}},T_{\text{HEMT}}$ are the added noise photons by JPA and linearised HEMT noise temperature respectively, $\text{SNR}_{\text{off}},\text{SNR}_{\text{on}}$ are the measured SNR when the pump is off and on respectively, $G$ is the gain of the JPA, and $L$ is the loss between the JPA and the HEMT. In our analysis, we assume that the $T_{\text{HEMT}}$ is equal to the nominal noise temperature in the datasheet of 3.63.63.6\,3.6K, and we estimate the loss between the IEJPA and the HEMT to be about 1.81.81.8\,1.8dB, mostly attributed to insertion loss for the circulators and the cable loss. The plot of estimated $n_{\text{add}}$ is shown by the orange curve in the middle panel of Fig. 2. We note that for a more precise characterization of the added noise, we need to use a shot-noise tunnel junction (SNTJ) noise thermometerSimoen et al. (2015), which was not available during this work. The green dotted line in Fig. 2(b) corresponds to the standard quantum limit of 0.50.50.5\,0.5 photons of noise added.
Figura 2: Measured (a) gain, (b) added noise, and (c) 1dB compression point for the IEJPA operated at a pump frequency and pump power of 5.3425.3425.342\,5.342GHz and $-88\,$dBm respectively. In the plot (b), the solid line represents the estimated $n_{\text{add}}$ while the dashed line represents the standard quantum limit of $0.5$ photons for a quantum limited amplifier.
We also measured the saturation power of the JPA (also called the $1$ dB compression point), which is the power of the input signal tone at which the gain decreases by $1$ dB. To carry out this measurement, we varied the power of the input signal tone while keeping the external magnetic flux, pump power, and pump frequency at the optimal values as mentioned above. The measured saturation power was around $-114$ dBm over the bandwidth of the JPA, as shown in Fig. 2(c).
As shown in Fig. 4, we find that our measured gain profile (blue, solid) is not explained sufficiently well by existing theory Roy et al. (2015), which accounts for only the lowest-order nonlinearity of the JPA connected to a linear transformer (red dashed). Therefore, we developed a model from first-principles by deriving the full Hamiltonian of the device, writing down the quantum Langevin equations and applying input-output theory to obtain the gain expression. Importantly, our theory retains the full sine nonlinearity of the SQUID of the JPA as well as the array of JJs in the transformer. We present a detailed description of our theory in the Supplementary Material. Here, we summarize the key steps of our modeling.
Figura 3: Equivalent circuit diagram of the device.
The equivalent circuit diagram of our device is shown in Fig. 3, where the impedance engineering is achieved by placing a series LC circuit (Transformer) between the transmission line (TL) and the JPA. In this diagram, $E_{J}$ and $C_{J}$ represent the Josephson energy and capacitance of the JPA, while $ME_{T}$ and $C_{T}$ correspond to the Josephson energy of each junction in the $M$-junction chain and capacitance of the transformer. The Hamiltonian for the device is
$\hat{H}$ $=\hat{H}_{J}+\hat{H}_{T}+\hat{H}_{\text{int}}+\hat{H}_{\text{env}}$
$=\hat{H}_{\text{sys}}+\hat{H}_{\text{int}}+\hat{H}_{\text{env}}.$ (2)
Here, $\hat{H}_{J}$ is the JPA Hamiltonian, which includes its full sine non-linearity. It is given by
$\hat{H}_{J}=-\frac{\hbar\Omega_{J}}{2}(\hat{A}_{J}-{\hat{A}_{J}}^{\dagger})^{2}-E_{J}\cos(c_{2}(\hat{A}_{J}+{\hat{A}_{J}}^{\dagger})),$ (3)
where $\Omega_{J}$ is the JPA resonance frequency, $\hat{A}_{J}$ is the destruction operator for the intracavity mode of the JPA, $c_{2}=\frac{2\pi}{\phi_{0}}\sqrt{\frac{\hbar Z_{J}}{2}}$ is the scaled zero-point phase fluctuation across the JPA junction in which $Z_{J}=\Omega_{J}L_{J}$ is the linearized impedance of the JPA mode with linearized inductance $L_{J}=\phi_{0}^{2}/4\pi^{2}E_{J}$. The transformer Hamiltonian $\hat{H}_{T}$ has the form
$\hat{H}_{T}$ $=-\frac{\hbar\Omega_{T,\text{eff}}}{2}(\hat{A}_{T}-{\hat{A}_{T}}^{\dagger})^{2}$
$-M^{2}E_{T}\cos(c_{1}\frac{(\hat{A}_{T}+{\hat{A}_{T}}^{\dagger})}{M}-c_{2}\frac{(\hat{A}_{J}+{\hat{A}_{J}}^{\dagger})}{M}),$ (4)
where the first term corresponds to the capacitor of the transformer modified by the environment and the second for an array of $M$ identical JJs. Here, $\hat{A}_{T}$ is the annihilation operator for the intracavity mode of the transformer, $c_{1}=\frac{2\pi}{\phi_{0}}\sqrt{\frac{\hbar Z_{T}}{2}}$ is the scaled zero-point phase fluctuation across the JJ chain in which $Z_{T}=\Omega_{T,\text{eff}}L_{T}$ is the transformer mode impedance with an effective linearized inductance $L_{T}=\phi_{0}^{2}/4\pi^{2}E_{T}$, and $c_{2}$ is as defined above.
The term $\hat{H}_{\text{int}}$ describes the interaction of the transformer with the environment and is given by
[…]
where $R$ is the impedance of the environment, and other quantities are as previously defined. In the above equation, we make an approximation by removing the explicit dependence on $\omega$ by replacing $\omega$ with $\Omega_{T,\text{eff}}$ as for the frequency ranges of interest, $\sqrt{\omega/\Omega_{T,\text{eff}}}\approx 1$. The interaction Hamiltonian can then be expressed as
[…]
where $\kappa=\frac{R\Omega_{T,\text{eff}}}{Z_{T}}$ is the linewidth of the transformer.
The Hamiltonian $\hat{H}_{\text{env}}$ describing the environment is
[…]
where $\hat{b}(\omega)$ is the annihilation operator for the TL at frequency $\omega$.
Starting from the Hamiltonian 2, we can write down the quantum Langevin equations for the amplitudes associated with the transformer and the JPA:
$\frac{d\hat{A}_{T}}{dt}=-\frac{i}{\hbar}[\hat{A}_{T},\hat{H}_{\text{sys}}]-\frac{\kappa}{2}\hat{A}_{T}+\sqrt{\kappa}\hat{A}_{\text{in}}$ (8)
$\frac{d\hat{A}_{J}}{dt}=-\frac{i}{\hbar}[\hat{A}_{J},\hat{H}_{\text{sys}}]$ (9)
In the presence of a strong classical pump and a weak quantum signal of interest, we can break the solution into two parts by considering the classical evolution of the modes under a strong drive followed by looking at weak quantum perturbations to the mode occupation. To this end, we represent the annihilation operators of the transformer and JPA modes as $\hat{A}_{T}(t)=(\alpha_{T}+\hat{a}_{T}(t))e^{-i\omega_{p}t}$ and $\hat{A}_{J}(t)=(\alpha_{J}+\hat{a}_{J}(t))e^{-i\omega_{p}t}$ where $\alpha_{T},\hat{a}_{T}$ and $\alpha_{J},\hat{a}_{J}$ are classical pump amplitudes and quantum fluctuations around them inside the transformer and JPA respectively and $\omega_{p}$ is the pump frequency.
First, setting the quantum fluctuations to zero ($\hat{a}_{T},\hat{a}_{J}\rightarrow 0$), we obtain the equations of motion for the classical pump inside the two modesKochetov and Fedorov (2015),
$\dot{\alpha}_{T}$ $=i\omega_{p}\alpha_{T}-\left(\frac{\kappa+i\Omega_{T,\text{eff}}}{2}\right)\alpha_{T}$
$+\sqrt{\kappa}\alpha_{\text{in}}-i\frac{E_{T}}{\hbar}c_{1}J_{1}(A_{\text{eff}})e^{i\phi_{\text{eff}}}$ (10)
$\dot{\alpha}_{J}$ $=i\omega_{p}\alpha_{J}-i\frac{\Omega_{J}}{2}\alpha_{J}$
$+i\frac{E_{T}}{\hbar}c_{2}J_{1}(A_{\text{eff}})e^{i\phi_{\text{eff}}}-i\frac{E_{J}}{\hbar}c_{2}J_{1}(A_{\text{jpa}})e^{i\phi_{\text{jpa}}}$ (11)
Here, $A_{\text{jpa}},A_{\text{eff}}$ and $\phi_{\text{jpa}}$, $\phi_{\text{eff}}$ represent the amplitudes and phases of the classical pumps across the JPA and the transformer Josephson junctions respectively and are defined as
$A_{\text{eff}}e^{i\phi_{\text{eff}}}=2(c_{1}\alpha_{T}-c_{2}\alpha_{J})/M$
$A_{\text{jpa}}e^{i\phi_{\text{jpa}}}=2(c_{2}\alpha_{J}).$
These equations for the pump can be integrated starting from empty cavities to obtain the correct steady-state solutions to the pump occupation in the modes.
Next, using the steady-state solutions for the pump occupations, the equations of motion for the weak signal in the cavity can be obtained by expanding the Langevin equations to first order in the perturbations $\hat{a}_{T},\hat{a}_{J}$ in the frame rotating at the pump frequecny $\omega_{p}$. In the frequency domain with the “signal” frequencies being $\omega_{s}$, defining $\Delta=\omega_{s}-\omega_{p}$, under a rotating wave approximation, the quantum equations of motion for the intracavity perturbations are
$-i\Delta\hat{a}_{T}(\Delta)$ $=\left(i\omega_{p}-i\frac{\Omega_{T,\text{eff}}}{2}-\frac{\kappa}{2}\right)\hat{a}_{T}(\Delta)+\sqrt{\kappa}\hat{a}_{\text{in}}(\Delta)$
$-i\frac{E_{T}}{\hbar}c_{1}J_{0}(A_{\text{eff}})(c_{1}\hat{a}_{T}(\Delta)-c_{2}\hat{a}_{J}(\Delta))$
$+i\frac{E_{T}}{\hbar}c_{1}J_{2}(A_{\text{eff}})\left(c_{1}\hat{a}_{T}^{\dagger}(-\Delta)-c_{2}\hat{a}_{J}^{\dagger}(-\Delta)\right)e^{i2\phi_{\text{eff}}},$ (12)
$-i\Delta\hat{a}_{J}(\Delta)$ $=i\left(\omega_{p}-\frac{\Omega_{J}}{2}\right)\hat{a}_{J}(\Delta)-i\frac{E_{J}}{\hbar}c_{2}^{2}J_{0}(A_{\text{jpa}})\hat{a}_{J}(\Delta)$
$+i\frac{E_{J}}{\hbar}c_{2}^{2}J_{2}(A_{\text{jpa}})\hat{a}^{\dagger}_{J}(-\Delta)e^{i2\phi_{\text{jpa}}}$
$+i\frac{E_{T}}{\hbar}c_{2}J_{0}(A_{\text{eff}})(c_{1}\hat{a}_{T}(\Delta)-c_{2}\hat{a}_{J}(\Delta))$
$-i\frac{E_{T}}{\hbar}c_{2}J_{2}(A_{\text{eff}})\left(c_{1}\hat{a}^{\dagger}_{T}(-\Delta)-c_{2}\hat{a}^{\dagger}_{J}(-\Delta)\right)e^{i2\phi_{\text{eff}}}.$ (13)
The coupled equations for $\hat{a}_{T}(\Delta)$ and $\hat{a}_{J}(\Delta)$ along with their Hermitian conjugates, and the input-output relation $\sqrt{\kappa}\hat{a}_{T}(\Delta)=\hat{a}_{\text{in}}(\Delta)+\hat{a}_{\text{out}}(\Delta)$, can be used to derive a scattering relation of the form
$\hat{a}_{\text{out}}(\Delta)=\sqrt{G(\Delta)}\hat{a}_{\text{in}}(\Delta)+\sqrt{G(\Delta)-1}{\hat{a}_{\text{in}}}^{\dagger}(-\Delta).$ (14)
In the above, $\hat{a}_{\text{in}}(\Delta)$ and ${\hat{a}_{\text{in}}}(-\Delta)$ represent the input operators for the signal and idler modes, respectively, while $\hat{a}_{\text{out}}(\Delta)$ is the output operator for the signal mode in the pump frame. Here, $G(\Delta)$ represents the gain of the signal mode and is the quantity we wish to compute from our theory model in order to compare to the experimentally measured gain profile.
Using our theory model, we have numerically computed the gain profile as follows. First, we solve Eq. 10 and Eq. 11 for the steady state solution starting with $\alpha_{T}(t=0)=\alpha_{J}(t=0)=0$ with a specified input pump power, $\alpha_{\text{in}}$ and a set of device parameters to obtain the long-time limit solutions $\alpha_{T}(t\rightarrow\infty),\alpha_{J}(t\rightarrow\infty)$. These steady-state pump occupations are then substituted in Eq. 12 and Eq. 13, which we use to calculate the gain. The device parameters used in the simulations are $\omega_{p}/{2\pi}=5.347\text{ GHz}$, $\Omega_{J}/{2\pi}=6.5\text{ GHz}$, $\Omega_{T,\text{eff}}/{2\pi}=6.218\text{ GHz}$, $L_{J}=0.37\text{ nH}$, and $L_{T}=1.9\text{ nH}$, where the relevant inductances are the sums of linearized inductance due to the junctions and geometric inductance due to the device geometry. The extra capacitance due to coupling to the environment was calculated for the experimental setup with a 444\,4GHz bandwidth, external impedance $R\sim 50\,\Omega$, and $C_{T}\sim 420\,\text{fF}$. In addition, the input pump power was set to align with the experimental conditions.
Figure 4 compares experimental results (blue, solid) with simulations for the IEJPA considering both finite and full nonlinearity of the Josephson junctions depicted as red dashed and green dotted lines respectively. The red theory plot considering only finite JPA nonlinearity does not match the experimental data. This observation motivated the inclusion of full junction non-linearity to better match the experimental results. The theoretical predictions considering full non-linearity predicts a reduction in the bandwidth over which appreciable gain is achievable, and hence partially explains the reduced bandwidth observed in the experiment. However, it does not match the measured profile well at frequencies far from the central frequency ($\omega_{p}$). This discrepancy may arise from the Markov approximation used in transitioning from Eq.5 to Eq.6, where $\kappa$ is approximated as $\kappa=2\pi\lambda^{2}({\Omega_{T,\text{eff}}})$. This approximation holds true primarily for devices with a high Q factor, which is not the case with our device. For comparison, the JPA with full sine nonlinearity without engineering is also shown in the plot as an orange dash-dotted line, where the device parameters used in the simulation are $\omega_{p}/{2\pi}=5.347\text{ GHz}$, $\Omega_{J}/{2\pi}=7.05\text{ GHz}$, and $L_{J}=0.323\text{ nH}$. The bandwidth enhancement due to impedance engineering is clearly evident when comparing the experimentally measured gain profile to this curve.
Figura 4: Comparision of experimentally measured gain with different theoretical models of the IEJPA gain for $\omega_{p}=2\pi\times 5.347\,$GHz. Here, the blue solid line shows the experimentally measured gain. The non-solid lines all represent gain obtained from different models of the pumped IEJPA. The yellow dash-dotted line shows the gain obtained by considering the circuit in Fig. 3 without the transformer, but considering the full nonlinearity of the JPA junctions. The red dashed line represents the gain calculated for the IEJPA while considering only the quartic nonlinearities of the pumped junctions. Finally, the green dotted line shows the gain calculated for the pumped IEJPA while considering the full pumped junction nonlinearities up to the rotating wave approximation.
In summary, we have fabricated and characterized a broadband impedance-engineered JPA in a single-step fabrication procedure. We measured a gain of 181818\,18dB over a 400 MHz bandwidth with near quantum-limited noise performance and saturation power of -114dBm. Our experimental results are comparable to state-of-the-art values reported in literature - impedance engineering with gain 202020\,20dB, bandwidth 640640640\,640MHz and saturation power $-110\,$dBm Roy et al. (2015) and impedance matching with gain 151515\,15dB, bandwidth 700700700\,700MHz and saturation power $-108\,$dBm Mutus et al. (2014) - while using a much simpler device fabrication process. The simpler fabrication process enables the faster and more reliable realisation of near-quantum-limited broadband amplifiers for multi-qubit readout Vijay (2020) or other sensing applications. Further extensions to this work can involve improving the saturation power by using an array of SQUIDs in the JPA and increasing the number of junctions in the transformer array to decrease transformer non-linearity. Furthermore, similar devices can be designed which use three-wave (instead of four-wave) mixing, which due to the large frequency separation of the pump and signal allow the pump tone to be easily filtered without affecting the amplified signal. Such devices can hence be used as a broadband squeezed microwave source for high-precision measurements where noise in one quadrature is below vacuum noiseMehmet et al. (2010). The compact footprint of the devices makes its on-chip integration into a qubit architecture possible, thereby mitigating insertion loss due to chip-to-chip interconnects Eddins et al. (2019).
The authors acknowledge the generous support of the Space Technology Cell at IISc and ISRO through the project STC-0444(2022). The authors also acknowledge the generous support of the Ministry of Electronics and Information Technology of the Govt. of India, under the centre of Excellence of Quantum Technology at the Indian Institute of Science, as well as the office of Principle Scientific Advisor, Govt. of India. SH and AP acknowledge the support of the Kishore Vaigyanik Protsahan Yojana (KVPY). AS acknowledges the support of a New Faculty Initiation Grant (NFIG) from IIT Madras.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Riferimenti bibliografici
  • Yurke et al. (1989) B. Yurke, L. Corruccini, P. Kaminsky, L. Rupp, A. Smith, A. Silver, R. Simon,  and E. Whittaker, “Observation of parametric amplification and deamplification in a josephson parametric amplifier,” Physical Review A 39, 2519 (1989).
  • Vijay, Slichter, and Siddiqi (2011) R. Vijay, D. H. Slichter,  and I. Siddiqi, “Observation of quantum jumps in a superconducting artificial atom,” Phys. Rev. Lett. 106, 110502 (2011).
  • Hatridge et al. (2013) M. Hatridge, S. Shankar, M. Mirrahimi, F. Schackert, K. Geerlings, T. Brecht, K. M. Sliwa, B. Abdo, L. Frunzio, S. M. Girvin, R. J. Schoelkopf,  and M. H. Devoret, “Quantum back-action of an individual variable-strength measurement,” Science 339, 178–181 (2013)https://www.science.org/doi/pdf/10.1126/science.1226897 .
  • Lin et al. (2013) Z. Lin, K. Inomata, W. Oliver, K. Koshino, Y. Nakamura, J. Tsai,  and T. Yamamoto, “Single-shot readout of a superconducting flux qubit with a flux-driven josephson parametric amplifier,” Applied Physics Letters 103, 132602 (2013).
  • Roy and Devoret (2016) A. Roy and M. Devoret, “Introduction to parametric amplification of quantum signals with josephson circuits,” Comptes Rendus Physique 17, 740–755 (2016), quantum microwaves / Micro-ondes quantiques.
  • Caves (1982) C. M. Caves, “Quantum limits on noise in linear amplifiers,” Phys. Rev. D 26, 1817–1839 (1982).
  • Krantz et al. (2016) P. Krantz, A. Bengtsson, M. Simoen, S. Gustavsson, V. Shumeiko, W. D. Oliver, C. M. Wilson, P. Delsing,  and J. Bylander, “Single-shot read-out of a superconducting qubit using a Josephson parametric oscillator,” Nature Communications 7, 11417 (2016).
  • Mallet et al. (2011) F. Mallet, M. Castellanos-Beltran, H. Ku, S. Glancy, E. Knill, K. Irwin, G. Hilton, L. Vale,  and K. Lehnert, “Quantum state tomography of an itinerant squeezed microwave field,” Physical Review Letters 106, 220502 (2011).
  • Yurke et al. (1988) B. Yurke, P. G. Kaminsky, R. E. Miller, E. A. Whittaker, A. D. Smith, A. H. Silver,  and R. W. Simon, “Observation of 4.2-k equilibrium-noise squeezing via a josephson-parametric amplifier,” Phys. Rev. Lett. 60, 764–767 (1988).
  • Eichler et al. (2011) C. Eichler, D. Bozyigit, C. Lang, M. Baur, L. Steffen, J. Fink, S. Filipp,  and A. Wallraff, “Observation of two-mode squeezing in the microwave frequency domain,” Physical Review Letters 107, 113601 (2011).
  • Castellanos-Beltran and Lehnert (2007) M. Castellanos-Beltran and K. Lehnert, “Widely tunable parametric amplifier based on a superconducting quantum interference device array resonator,” Applied Physics Letters 91, 083509 (2007).
  • Yamamoto et al. (2008) T. Yamamoto, K. Inomata, M. Watanabe, K. Matsuba, T. Miyazaki, W. D. Oliver, Y. Nakamura,  and J. S. Tsai, “Flux-driven Josephson parametric amplifier,” Applied Physics Letters 93 (2008), 10.1063/1.2964182, 042510, https://pubs.aip.org/aip/apl/article-pdf/doi/10.1063/1.2964182/14123646/042510_1_online.pdf .
  • Mutus et al. (2013) J. Mutus, T. White, E. Jeffrey, D. Sank, R. Barends, J. Bochmann, Y. Chen, Z. Chen, B. Chiaro, A. Dunsworth, et al., “Design and characterization of a lumped element single-ended superconducting microwave parametric amplifier with on-chip flux bias line,” Applied Physics Letters 103, 122602 (2013).
  • Zhou et al. (2014) X. Zhou, V. Schmitt, P. Bertet, D. Vion, W. Wustmann, V. Shumeiko,  and D. Esteve, “High-gain weakly nonlinear flux-modulated josephson parametric amplifier using a squid array,” Phys. Rev. B 89, 214517 (2014).
  • Vijay (2020) R. Vijay, “Multiplexed readout of four qubits in 3d cqed architecture using broadband jpa,” Bulletin of the American Physical Society 65 (2020).
  • CULLEN (1958) A. L. CULLEN, “A Travelling-Wave Parametric Amplifier,” Nature 181, 332–332 (1958).
  • Feldman, Parrish, and Chiao (1975) M. J. Feldman, P. T. Parrish,  and R. Y. Chiao, “Parametric amplification by unbiased Josephson junctions,” Journal of Applied Physics 46, 4031–4042 (1975)https://pubs.aip.org/aip/jap/article-pdf/46/9/4031/7955853/4031_1_online.pdf .
  • Wahlsten, Rudner, and Claeson (1977) S. Wahlsten, S. Rudner,  and T. Claeson, “Parametric amplification in arrays of Josephson tunnel junctions,” Applied Physics Letters 30, 298–300 (1977)https://pubs.aip.org/aip/apl/article-pdf/30/6/298/7738589/298_1_online.pdf .
  • Ho Eom et al. (2012) B. Ho Eom, P. K. Day, H. G. LeDuc,  and J. Zmuidzinas, “A wideband, low-noise superconducting amplifier with high dynamic range,” Nature Physics 8, 623–627 (2012).
  • Yaakobi et al. (2013) O. Yaakobi, L. Friedland, C. Macklin,  and I. Siddiqi, “Parametric amplification in josephson junction embedded transmission lines,” Phys. Rev. B 87, 144301 (2013).
  • Zorin (2019) A. Zorin, “Flux-driven josephson traveling-wave parametric amplifier,” Phys. Rev. Appl. 12, 044051 (2019).
  • Macklin et al. (2015) C. Macklin, K. O’brien, D. Hover, M. Schwartz, V. Bolkhovsky, X. Zhang, W. Oliver,  and I. Siddiqi, “A near–quantum-limited josephson traveling-wave parametric amplifier,” Science 350, 307–310 (2015).
  • White et al. (2015) T. White, J. Mutus, I.-C. Hoi, R. Barends, B. Campbell, Y. Chen, Z. Chen, B. Chiaro, A. Dunsworth, E. Jeffrey, et al., “Traveling wave parametric amplifier with josephson junctions using minimal resonator phase matching,” Applied Physics Letters 106, 242601 (2015).
  • (24) D.M.Pozar, “Microwave engineering (john wiley & sons, 2019),”  .
  • Lu et al. (2022) Y. Lu, W. Xu, Q. Zuo, J. Pan, X. Wei, J. Jiang, Z. Li, K. Zhang, T. Guo, S. Wang, et al., “Broadband josephson parametric amplifier using lumped-element transmission line impedance matching architecture,” Applied Physics Letters 120, 082601 (2022).
  • Mutus et al. (2014) J. Y. Mutus, T. C. White, R. Barends, Y. Chen, Z. Chen, B. Chiaro, A. Dunsworth, E. Jeffrey, J. Kelly, A. Megrant, et al., “Strong environmental coupling in a josephson parametric amplifier,” Applied Physics Letters 104, 263513 (2014).
  • Roy et al. (2015) T. Roy, S. Kundu, M. Chand, A. Vadiraj, A. Ranadive, N. Nehra, M. P. Patankar, J. Aumentado, A. Clerk,  and R. Vijay, “Broadband parametric amplification with impedance engineering: Beyond the gain-bandwidth product,” Applied Physics Letters 107, 262601 (2015).
  • Grebel et al. (2021) J. Grebel, A. Bienfait, É. Dumur, H.-S. Chang, M.-H. Chou, C. Conner, G. Peairs, R. Povey, Y. Zhong,  and A. Cleland, “Flux-pumped impedance-engineered broadband josephson parametric amplifier,” Applied Physics Letters 118, 142601 (2021).
  • Simoen et al. (2015) M. Simoen, C. Chang, P. Krantz, J. Bylander, W. Wustmann, V. Shumeiko, P. Delsing,  and C. Wilson, “Characterization of a multimode coplanar waveguide parametric amplifier,” Journal of Applied Physics 118, 154501 (2015).
  • Kochetov and Fedorov (2015) B. A. Kochetov and A. Fedorov, “Higher-order nonlinear effects in a josephson parametric amplifier,” Phys. Rev. B 92, 224304 (2015).
  • Mehmet et al. (2010) M. Mehmet, H. Vahlbruch, N. Lastzka, K. Danzmann,  and R. Schnabel, “Observation of squeezed states with strong photon-number oscillations,” Physical Review A 81, 013814 (2010).
  • Eddins et al. (2019) A. Eddins, J. Kreikebaum, D. Toyli, E. Levenson-Falk, A. Dove, W. Livingston, B. Levitan, L. Govia, A. Clerk,  and I. Siddiqi, “High-efficiency measurement of an artificial atom embedded in a parametric amplifier,” Physical Review X 9, 011004 (2019).
See pages 1, of Figures/supplementary.pdf See pages 2, of Figures/supplementary.pdf See pages 3, of Figures/supplementary.pdf See pages 4, of Figures/supplementary.pdf See pages 5, of Figures/supplementary.pdf See pages 6, of Figures/supplementary.pdf See pages 7, of Figures/supplementary.pdf See pages 8, of Figures/supplementary.pdf See pages 9, of Figures/supplementary.pdf See pages 10, of Figures/supplementary.pdf See pages 11, of Figures/supplementary.pdf